OpenStudy (anonymous):

Solve 6x - 2 = 4x + 13 for x.

OpenStudy (anonymous):

6x - 4x = 13 + 2 2x = 15 x = 7.5

OpenStudy (anonymous):

what is 7.5 as a fraction

OpenStudy (anonymous):

3/12

OpenStudy (anonymous):

///

OpenStudy (anonymous):

\[7\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

???

OpenStudy (anonymous):

my chooses are 15/2 11/2 3/12

OpenStudy (anonymous):

ah, 15/2 then

OpenStudy (anonymous):

15 / 2 = 7.5 as any calculator will tell you....

OpenStudy (anonymous):

thank you are u sure that's it

OpenStudy (anonymous):

I'm 500/5 % sure